The quantity supplied of a commodity at a price of Birr 10 per unit is 500 units. Its price elasticity of supply is 0.5 A) Calculate the price at which its quantity supplied is 600 units. B) What quantity of this commodity will the seller supplied when the price rise at Birr 12 per unit? C) What quantity of this commodity will the sellers supply when price rises by Birr 4 per unit?


1
Expert's answer
2022-01-12T18:09:17-0500

A) The price at which its quantity supplied is 600 units is:

"P = 10*(1 + 100\/500*2) = 14."

B) The quantity of this commodity the seller will supply when the price rise at Birr 12 per unit is:

"Qs = 500*(1 + 2\/10*0.5) = 550" units. 

C) The quantity of this commodity that the sellers will supply when price rises by Birr 4 per unit is 600 units.